### v1

@@ Abstract
Altermagnets are a novel platform to realize exotic electromagnetic properties distinct from those of conventional ferromagnets and antiferromagnets. We report results of micro-focused angle-resolved photoemission spectroscopy (ARPES) on RuO2, of which altermagnetic nature has been under fierce debate. We have elucidated the band structure of the (100), (110) and (101) surfaces of a bulk single crystal. We found that, irrespective of the surface orientation, the experimental band structures obtained by ARPES commonly show a semi-quantitative agreement with the bulk-band calculation for the nonmagnetic phase, but display a severe disagreement with that for the antiferromagnetic phase. Moreover, spin-resolved ARPES signifies a negligible spin polarization for the bulk bands. These results suggest the absence of antiferromagnetism and altermagnetic spin splitting. Furthermore, we identified a flat surface band and a dispersive one near the Fermi level at the (100)/(110) and (101) surfaces, respectively, both of which are attributed to the topological surface states associated with the bulk Dirac nodal lines. The present ARPES results suggest that the crystal-orientation-dependent topological surface/interface states need to be taken into account to properly capture the transport and catalytic characteristics of RuO2.
@@ I INTRODUCTION
Ru with rutile structure [see Fig. 1(a)] has attracted significant attention because of its outstanding properties associated with the spin-charge coupling [12, 15, 21, 22], provoking active discussions on whether these properties are attributable to the altermagnetic spin splitting [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In fact, the observed magnitude of the spin Hall effect is two orders of magnitude smaller than the expectation. When one assumes a collinear antiferromagnetic (AFM) structure [12, 32, 33, 34], the oxygen atoms surrounding the Ru atoms break the symmetry [Fig. 1(a)], resulting in a -wave-like anisotropic spin splitting of maximally 0.8 eV along the [101] direction with gap nodes in the (100) and (010) planes [15, 22]. This anisotropic splitting was theoretically predicted to give rise to a giant anomalous Hall effect and highly efficient spin-charge conversion, both of which show a strong crystallographic-axis dependence [15, 22, 21]. These predictions have been examined experimentally using thin-film samples with various crystal orientations [22, 23, 24, 25, 27, 28, 29, 30], as highlighted by the observation of anomalous Hall conductivity comparable to that of Fe in a (110)-oriented sample [22] and high spin-charge conversion efficiency in a (101)-oriented sample [21, 25]. RuO2 also shows intriguing physical and chemical properties such as anisotropic anomalous Nernst effect [35], anisotropic tunneling magnetoresistance [36, 37], terahertz emission [38, 39], and catalytic reactivity on the (110) surface [40], while their relationship with the electronic states has yet to be clarified.
@@
While the experimental evidence showing the crystal-orientation-dependence to support the anisotropic spin splitting is accumulating in RuO2 [22, 24, 23, 26, 27, 28, 29, 30], there exist some contradictory reports from the neutron scattering and SR [41, 42] that do not support the magnetic order. Consequently, there remain two key unresolved questions; i.e. (i) whether RuO2 truly exhibits altermagnetism, and (ii) what microscopic electronic structures underlie the observed strong crystal-orientation-dependent transport properties. Although angle-resolved photoemission spectroscopy (ARPES) is considered to be one of essential techniques for addressing these issues through direct observation of the band structure, current ARPES studies are limited to the (110) crystal plane, leaving a comprehensive understanding of its electronic structure incomplete [43, 44, 45, 46]. Moreover, even within ARPES studies on (110)-oriented single crystals, the results remain highly controversial, with conflicting reports between the paramagnetic (PM) band degeneracy and the altermagnetic band splitting [46, 45]. It is thus highly desirable to carry out the comprehensive ARPES measurements on different surface orientations to clarify the universality and dissimilarity in the surface-dependent electronic states.
@@
In this article, we report a spin-resolved ARPES study with a RuO2 bulk single crystal which has three different surface planes of (100), (110) and (101), in combination with first-principles calculations, to address the key questions above. We show that the absence of altermagnetic spin splitting is a universal feature irrespective of crystal orientations. We also found energy bands crossing the Fermi level () associated with the surface states (SS) in the (, ) region where no corresponding bands exist in the bulk-band calculation. The effective mass of SS strongly depends on the crystal orientation. We discuss implications of the present results in relation to the topological properties.
@@ III Electronic structure from (100), (110), and (101) surfaces
Although previous ARPES measurements with bulk single crystals and thin films were carried out on the (110) surface [43, 44, 45, 46], we found that the bulk crystal can be cleaved nicely also on the (100) and (101) surfaces [indicated by the shaded area in Fig. 1(a)]. These three types of flat surfaces are identified as facets of an as-grown single crystal as shown in Fig. 1(b). We have characterized the orientation of these surfaces by x-ray Laue backscattering measurements [Fig. 1(c)] in which clear and sharp diffraction spots, consistent with the symmetry of each surface, are well recognized.
@@
Figures 2(b) and 2(c) show the experimental FSs for the (100) surface plotted against the in-plane wave vectors ( and ) obtained for the plane, which are directly compared with the slice of bulk FSs obtained with the NM and AFM calculations, respectively. The ARPES intensity signifies some characteristic features, such as the absence of spectral weight around the R point and complicated intensity distribution around the point. These features have a periodicity of bulk BZ with two-fold symmetry, indicative of the successful cleavage at the (100) surface. As shown in Fig. 2(b), the calculated FSs in the NM phase show an excellent agreement with the ARPES intensity. In particular, rounded triangular pockets (called ) around the Z point intersecting along the ZR cut and the absence of FS around the R point seen in the experiment are nicely reproduced in the NM calculation. On the other hand, as shown in Fig. 2(c), the calculated FS in the AFM phase is too simple; it just contains small pockets centered at the and Z points, in contradiction to the ARPES result. We have confirmed by systematically changing the on-site coulomb energy in the AFM calculation that the discrepancy between the experiment and AFM calculation is not attributed to the incorrect choice of (for details, see Supplementary Note 1).
@@
We identified an experimental feature which is not reproduced in neither the NM- nor AFM-calculations. That is a nearly flat band in the vicinity of along the XR and Z cuts in Fig. 2(h) [orange circles in Figs. 2(e), 2(g), and 2(i)]. It is assigned to the SS, as supported by its -independent band dispersion and also by our surface-weight calculations (for details, see Supplementary Notes 2 and 3). In fact, the invariance at the point can be inferred by looking at the commonly identified near- flat band at X, , and points in Figs. 2(d), 2(f) and 2(h) all of which are projected onto the point in the surface Brillouin zone (BZ), despite a strong intensity modulation due to the matrix-element effect of photoelectron intensity.
@@
The argument above is further corroborated by a comparison of band dispersions at between ARPES and calculations in Figs. 2(d) and 2(e). Besides the B1 band, the ARPES intensity along the MA cut signifies another band called B2 which shows Dirac-like crossing with the B1 band around . This crossing is nicely reproduced by the NM calculation in the similar (, ) region [Fig. 2(e)], and is attributed to the bulk Dirac nodal line (DNL) with the quadruple band degeneracy appearing commonly in rutile oxides [43, 56, 57, 58]. Interestingly, the SS shows up from this DNL and resides on one side (left-hand side) of the DNL along the MA cut; we will come back to this point later. Although the AFM calculation also shows a Dirac-crossing behavior at (open red diamond), its point and the energy dispersion corresponding to the B2 band are very different, again supporting the agreement with the NM calculation. As can be seen in Fig. 2(g), the AFM calculation shows the altermagnetic spin splitting at = 0.5 (the off-nodal plane), leading to the doubling of some bands with respect to the NM case. The ARPES intensity in Fig. 2(f) exhibits no clear signature of such band doubling, as inferred from the smooth evolution of spectral features from = 0 [Fig. 2(h)] to 1.0 [Fig. 2(d)] through = 0.5 [Fig. 2(f)]. We carried out spin-resolved ARPES measurements at points where the altermagnetic band splitting is theoretically predicted [red line in Fig. 2(f)], but found that the spin polarization for the bulk bands always keeps zero within experimental uncertainty (for details, see Supplementary Note 4). All these results indicate that the electronic structure of the (100) surface does not support the altermagnetic band splitting.
@@
We found that the ARPES data for the (110) surface [see Fig. 3(a)] share several common features with the (100) surface. The FS mapping against the in-plane wave vectors in Fig. 3(b) obtained in the plane [ZAM plane; red-shaded plane in Fig. 3(a)] signifies a two-fold-symmetric intensity pattern with the periodicity obeying that of the bulk BZ, consistent with the previous reports [43, 44, 46, 45]. A side-by-side comparison of the FS mapping with the NM and AFM calulcations in Figs. 3(b) and 3(c), respectively, signifies that the overall FS shape surrounding the point is reproduced well by the NM calculation whereas the size of -centered FS in the AFM calculation is much smaller than that of the experiment. The ARPES intensity shown in Fig. 3(d) along the ZAM cut signifies several dispersive bands. Comparison of the ARPES-derived band dispersion with the calculated band dispersions for the NM and AFM cases in Fig. 3(e) reveals that the (, ) position of experimental bands are semi-quantitatively reproduced in the NM calculation, whereas the AFM calculation apparently shows a mismatch in the energy position. Similarly to the case of the (100) surface, we found evidence for a flat SS around the point with its surface nature suggested from the -invariance of the peak position as shown in Fig. 3(f). It is remarked here that the previous ARPES study supporting the altermagneitc band splitting [46] attributed this flat band to the flat band at 0.4 eV seen in the AFM calculation along the M cut (red dashed line) by largely shifting the calculated position downward, while this amount of energy shift is unrealistic taking into account of nearly stoichiometric nature of the RuO2 single crystal.
@@
To clarify the band character, we show in Figs. 4(d) and 4(f) the ARPES intensity along representative cuts in the 0.0 and 1.0 planes, respectively, compared with the corresponding band dispersions obtained by the NM and AFM calculations in Figs. 4(e) and 4(g). One can recognize in Figs. 4(d) and 4(f) several dispersive bands, some of which cross to contribute to the pockets seen in Figs. 4(b) and 4(c). Although the distinction of more appropriate calculation (NM or AFM) is difficult compared to other surface planes, the experimental broad feature at 0.3–0.6 eV along the RAR cut in Fig. 4(f) appears to show a better agreement with the NM calculation in Fig. 4(g). Besides these bulk-originated features, one can find in both Figs. 4(d) and 4(f) several dispersive features within eV of , while they have no counterparts in the calculation. For example, an inner electron band that produces the hexagonal pocket in Fig. 4(b) appears in the (, ) region where the calculated bands in the NM phase are absent. This band displays no band dispersion along the out-of-plane wave vector, [Figs. 4(d) and 4(f)], supporting its surface origin. One can also recognize a hole band producing a small pocket at the X point associated with the pocket in Fig. 4(b). This band is not predicted by the NM calculation and has no dispersion along in the experiment [Fig. 4(g)], and hence, it is also attributed to the SS. Therefore, the (101) surface of RuO2 hosts several dispersive SS that contribute to the formation of surface-derived FS. It is noted here that the inclusion of SOC in the NM calculation slightly improves the matching between the experiment and calculation (see Supplementary Note 5), but a fatal disagreement is still found in the AFM calculation even when the SOC is included.
@@ IV Topological surface states
Here we summarize our key findings to discuss the implications of present results. We have succeeded in experimentally clarifying the band structure and FS of RuO2 for multiple crystal orientations, and found that the experimental band structure always shows a better agreement with the NM calculation compared to the AFM calculation irrespective of the crystal orientations. Furthermore, our spin-ARPES measurements definitively show no evidence for the spin polarization of the bulk bands. The present study supports the absence of antiferromagnetism and altermagnetic band splitting in bulk single crystal of RuO2, settling the fierce debates on the altermagnetic nature of RuO2, at least for the bulk crystal (note that, at the moment, we do not exclude the possibility that a RuO2 thin film shows altermagnetic behavior, because the crystal defects and lattice strain are not the same as those of bulk crystal). We also found that the existence of SS is a common characteristic of RuO2 irrespective of the crystal orientation, while its energy dispersion on the (101) surface shows dispersive characteristics distinct from those of the (100) and (110) surfaces hosting nearly flat SS around .
We propose that the key characteristics of SS observed at the three different crystal orientations are commonly explained in terms of the existence of DNL in bulk bands and associated topological SS. First-principles calculations for rutile oxides (RuO2, IrO2 and OsO2) [56] predicted the existence of bulk DNL accompanying topological SS [43, 57, 58], which is located at around a specific plane (2/3 in the unit of ; called here the DNL plane) and extends along the and axes as shown in Fig. 3(a). Actually, the SOC lifts the degeneracy of Dirac node by 0.1 eV and contributes to the spin Berry curvature [56]; see also Figs. S5(a) and S5(b). Being associated with this DNL, a large volume of region in bulk BZ (region with ; indicated by red dashed box) including the , X, and M points becomes a band-inverted region, whereas the remaining smaller region with including the Z, A, and R points is not inverted. Since the (100) and (110) planes vertically intersect the DNL plane, one can distinguish between inverted and non-inverted regions in the projected surface BZ. Namely, the topological SS is expected to emerge in the cut which crosses the surface projection of bulk , X, and M points, whereas it is absent in the cut connecting the surface projection of the bulk Z, R, and A lines.
We found that the observed SS exactly follows the above trend. Specifically, at the (100) surface, the SS appears along the XM, MA, and XR cuts, but not along the AR cut as seen in Fig. 2(d). Similarly, it appears along the X, XR, and Z cuts, but not along the ZR cut as seen in Fig. 2(h). At the (110) surface [Fig. 3(d)], the SS appears along the Z, AM, and M cuts, but not along the ZA cut. These arguments are also supported by the FS mapping in Figs. 2(a) and 3(b) where the overall intensity in the band-inverted area of is relatively enhanced due to the presence of SS. This is also corroborated by slab calculations detailed in Supplementary Note 3.
In contrast to the (100) and (110) planes, the (101) plane intersects the DNL plane with , so that the surface projection of inverted region covers the whole region of the (101) surface BZ. This situation makes it difficult to predict the actual location of SS and is also likely responsible for a more dispersive nature of the SS in the (101) surface compared to the (100) and (110) planes in which relationship between the (, ) locations of the DNL and associated SS is more straightforward. We thus infer that the SS appearing ubiquitously in all the three crystal orientations is a reflection of its topological nature associated with the bulk-band inversion in a wide 3D space, and strongly affects the surface characteristics of RuO2 such as spectroscopic, transport, and chemical properties. In this regard, it is worthwhile to comment on the possibility that the catalytic reaction particularly known for the (110) surface [40] may be promoted by the existence of topological SS [43], because high surface conductivity and robustness of the topological SS would be useful for the catalytic reaction [59, 60], working as a “topological catalyst”. In RuO2, the proximity of the topological SS to [ 20 meV ( 200 K) of ; see Fig. 3(f)] for the (110) and (100) surfaces would promote catalytic reaction at room temperature.
@@
While the present study focuses on the surface of bulk crystal, it would be natural to extend our arguments to the interface between RuO2 (including RuO2 thin film) and other materials. The present study suggests that the outstanding transport properties of RuO2 thin film such as the spin-charge conversion (spin splitting torque) and other physical properties that involve charge/spin transport across the interface should be discussed with taking into account the existence of prominent topological interface states besides the bulk DNL that strongly contributes to the spin Berry curvature [30, 56]. In other words, we caution that a theoretical model that simply deals with the altermagnetic splitting of bulk bands without incorporating the topological surface/interface states and bulk DNL is insufficient to fully account for some of the experimental results of RuO2.

### v2

@@ Abstract
Altermagnets are a novel platform to realize exotic electromagnetic properties distinct from those of conventional ferromagnets and antiferromagnets. We report results of micro-focused angle-resolved photoemission spectroscopy (ARPES) on RuO2, in which its altermagnetic nature has been under fierce debate in connection with crystal-orientation-dependent spintronic functionalities. By elucidating the band structure of the (100), (110) and (101) surfaces of a bulk single crystal using micro-ARPES, we found that, irrespective of the surface orientation, the experimental band structures show a good agreement with the bulk-band calculations for the nonmagnetic phase, but display a severe disagreement with those for the antiferromagnetic phase. Moreover, spin-resolved ARPES signifies a negligible spin polarization in the bulk bands, suggesting the absence of antiferromagnetism and altermagnetic spin splitting. In addition, we identified a nearly flat surface band and a dispersive one near the Fermi level at the (100)/(110) and (101) surfaces, respectively. Our first-principles calculations and analysis of Berry phase attribute these states to the topological surface bands emerging from the bulk Dirac nodal lines around the Fermi level. Our results indicate that such topological surface/interface states must be considered to understand the spintronic functionalities of RuO2, and may provide new insights into its catalytic characteristics.
@@ I INTRODUCTION
Ru with rutile structure [see Fig. 1(a)] has attracted significant attention because of its outstanding properties associated with the spin-charge coupling [3, 57, 20, 19], provoking active discussions on whether these properties are attributable to the altermagnetic spin splitting [19, 5, 33, 9, 6, 63, 65, 22, 66, 28]. When one assumes a collinear antiferromagnetic (AFM) structure [3, 7, 75, 21], the oxygen atoms surrounding the Ru atoms break the symmetry [Fig. 1(a)], resulting in a -wave-like anisotropic spin splitting of maximally 0.8 eV along the [101] direction with gap nodes in the (100) and (010) planes [57, 19]. This anisotropic splitting was theoretically predicted to give rise to a giant anomalous Hall effect and highly efficient spin-charge conversion, both of which show a strong crystallographic-axis dependence [57, 19, 20]. These predictions have been examined experimentally using thin-film samples with various crystal orientations [19, 5, 33, 9, 63, 65, 22, 66], as highlighted by the observation of anomalous Hall conductivity comparable to that of Fe in a (110)-oriented sample [19] and high spin-charge conversion efficiency in a (101)-oriented sample [20, 9]. RuO2 also shows intriguing physical and chemical properties such as anisotropic anomalous Nernst effect [74], anisotropic tunneling magnetoresistance [1, 53, 48], terahertz emission [42, 73], and catalytic reactivity on the (110) surface [49], while their relationship with the electronic states has yet to be clarified.
@@
While the experimental evidence showing the crystal-orientation-dependence to support the anisotropic spin splitting is accumulating in RuO2 [19, 33, 5, 6, 63, 65, 22, 66], there exist some contradictory reports from the neutron scattering and SR [24, 35] that do not support the magnetic order. Consequently, there remain two key unresolved questions; i.e. (i) whether RuO2 truly exhibits altermagnetism, and (ii) what microscopic electronic structures underlie the observed strong crystal-orientation-dependent transport properties. Although angle-resolved photoemission spectroscopy (ARPES) is considered to be one of essential techniques for addressing these issues through direct observation of the band structure, current ARPES studies are limited to the (110) crystal plane, leaving a comprehensive understanding of its electronic structure incomplete [30, 31, 29, 18, 41, 40]. Moreover, even within ARPES studies on (110)-oriented single crystals, the results remain highly controversial, with conflicting reports between the paramagnetic (PM) band degeneracy and the altermagnetic band splitting [40, 41] (for chronological overview of the ARPES studies on rutile oxides, see Supplemental Note 1[2]). This controversy cannot be resolved solely by the ARPES studies of the (110) orientation, because key spectral feature to distinguish the altermagnetic band splitting, a flat band near (which will be detailed later), may be attributed either to a spin-split bulk band [40] or to surface states [41]. Therefore, it is essential to comprehensively investigate the electronic states of multiple orientations to disentangle surface and bulk states as well as to clarify the universality and dissimilarity in the surface-dependent electronic states.
@@
In this article, we report a spin-resolved ARPES study with a RuO2 bulk single crystal which has three different surface planes of (100), (110) and (101), in combination with first-principles calculations, to address the key questions above. Our ARPES results provide strong spectroscopic evidence supporting the absence of altermagnetic spin splitting in bulk RuO2 irrespective of crystal orientations. We also found energy bands crossing the Fermi level () associated with the surface states (SS) in the (, ) region where no corresponding bands exist in the bulk-band calculation. The effective mass of SS strongly depends on the crystal orientation. We discuss implications of the present results in relation to the topological properties.
@@ III Electronic structure from (100), (110), and (101) surfaces
Although previous ARPES measurements with bulk single crystals and thin films were carried out on the (110) surface [30, 31, 29, 18, 41, 40], we found that the bulk crystal can be cleaved nicely also on the (100) and (101) surfaces [indicated by the shaded area in Fig. 1(a)]. These three types of flat surfaces are identified as facets of an as-grown single crystal as shown in Fig. 1(b). We have characterized the orientation of these surfaces by x-ray Laue backscattering measurements [Fig. 1(c)] in which clear and sharp diffraction spots, consistent with the symmetry of each surface, are well recognized.
@@
Figures 2(b) and 2(c) show the experimental FSs for the (100) surface plotted against the in-plane wave vectors ( and ) obtained for the plane, which are directly compared with the slice of bulk FSs obtained with the NM and AFM calculations, respectively. The ARPES intensity signifies some characteristic features, such as the absence of spectral weight around the R point and complicated intensity distribution around the point. These features have a periodicity of bulk BZ with two-fold symmetry, indicative of the successful cleavage at the (100) surface. As shown in Fig. 2(b), the calculated FSs in the NM phase show an excellent agreement with the ARPES intensity. In particular, rounded triangular pockets (called ) around the Z point intersecting along the ZR cut and the absence of FS around the R point seen in the experiment are nicely reproduced in the NM calculation. On the other hand, as shown in Fig. 2(c), the calculated FS in the AFM phase is too simple; it just contains small pockets centered at the and Z points, in contradiction to the ARPES result. We have confirmed by systematically changing the on-site coulomb energy in the AFM calculation that the discrepancy between the experiment and AFM calculation is not attributed to the incorrect choice of (for details, see Appendix A).
@@
We identified an experimental feature which is not reproduced in neither the NM- nor AFM-calculations. That is a nearly flat band in the vicinity of along the XR and Z cuts in Fig. 2(h) [near- region of orange circles in Figs. 2(e), 2(g), and 2(i)]. It is assigned to the surface state (SS), as supported by its -independent band dispersion and also by our slab calculations (for details, see Appendix B and F). In fact, the invariance at the point can be inferred by looking at the commonly identified near- flat band at X, , and points in Figs. 2(d), 2(f) and 2(h) all of which are projected onto the point in the surface Brillouin zone (BZ), despite a strong intensity modulation due to the matrix-element effect of photoelectron intensity. A clear identification of the sharp SS supports good surface quality. Moreover, the crystals used in this study is of substantially higher quality, with RRR of 400, than those in other recent reports in 2019–2025, 10 to 200 [51, 52, 27, 50, 21, 19, 33]. Previous ARPES reports did not mention the sample quality in terms of RRR as summarized in the table S1 in Supplemental Note 1[2].
@@
The argument above is further corroborated by a comparison of band dispersions at between ARPES and calculations in Figs. 2(d) and 2(e). Besides the B1 band, the ARPES intensity along the MA cut signifies another band called B2 which shows Dirac-like crossing with the B1 band around . This crossing is nicely reproduced by the NM calculation in the similar (, ) region [Fig. 2(e)], and is attributed to the bulk Dirac nodal line 1 (DNL1) with the quadruple band degeneracy appearing commonly in rutile oxides [30, 31, 61, 68, 47]. Interestingly, the SS shows the dispersion starting from this DNL1 along the MA cut; we will come back to this point later. Although the AFM calculation also shows a Dirac-crossing behavior at (open red diamond), its point and the energy dispersion corresponding to the B2 band are very different, again supporting the agreement with the NM calculation. As can be seen in Fig. 2(g), the AFM calculation shows the altermagnetic spin splitting at = 0.5 (the off-nodal plane), leading to the doubling of some bands with respect to the NM case. The ARPES intensity in Fig. 2(f) exhibits no clear signature of such band doubling, as inferred from the smooth evolution of spectral features from = 0 [Fig. 2(h)] to 1.0 [Fig. 2(d)] through = 0.5 [Fig. 2(f)]. We carried out spin-resolved ARPES measurements at points where the altermagnetic band splitting is theoretically predicted [red line in Fig. 2(f)], but found that the spin polarization for the bulk bands always keeps zero within experimental uncertainty (for details, see Appendix C). All these results indicate that the electronic structure of the (100) surface does not support the altermagnetic band splitting.
@@
We found that the ARPES data for the (110) surface [see Fig. 3(a)] share several common features with the (100) surface. The FS mapping against the in-plane wave vectors in Fig. 3(b) obtained in the plane [ZAM plane; red-shaded plane in Fig. 3(a)] signifies a two-fold-symmetric intensity pattern with the periodicity obeying that of the bulk BZ, consistent with the previous reports [30, 31, 29, 18, 40, 41]. A side-by-side comparison of the FS mapping with the NM and AFM calculations in Figs. 3(b) and 3(c), respectively, signifies that the overall FS shape surrounding the point is reproduced well by the NM calculation whereas the size of -centered FS in the AFM calculation is much smaller than that of the experiment. The ARPES intensity shown in Fig. 3(d) along the ZAM cut signifies several dispersive bands. Comparison of the ARPES-derived band dispersion with the calculated band dispersions for the NM and AFM cases in Fig. 3(e) reveals that the (, ) position of experimental bands are semi-quantitatively reproduced in the NM calculation, whereas the AFM calculation apparently shows a mismatch in the energy position. Similarly to the case of the (100) surface, we found evidence for a flat SS around the point with its surface nature suggested from the -invariance of the peak position as shown in Fig. 3(f). It is remarked here that the previous ARPES study supporting the altermagnetic band splitting [40] attributed this flat band to the flat band at 0.4 eV seen in the AFM calculation along the M cut (red dashed line) by largely shifting the calculated position downward, while this amount of energy shift is unrealistic taking into account of nearly stoichiometric nature of the RuO2 single crystal. It is emphasized here that ARPES measurements on different surface orientations play a crucial role in testing the validity of band assignments, since the surface and bulk band dispersions inherently exhibit different sensitivities to change in surface orientation. Moreover, transport data in RuO2 thin films are known to depend strongly on surface orientation (e.g. refs. [57, 59, 19, 33, 5, 6, 63, 65, 22, 66, 9, 20]). Because such transport properties reflect bulk electronic states, clarifying the variations in bulk-band dispersions across surface orientations is highly valuable for identifying the origin of the orientation-dependent transport properties.
@@
To clarify the band character, we show in Figs. 4(d) and 4(f) the ARPES intensity along representative cuts in the 0.0 and 1.0 planes, respectively, compared with the corresponding band dispersions obtained by the NM and AFM calculations in Figs. 4(e) and 4(g). One can recognize in Figs. 4(d) and 4(f) several dispersive bands, some of which cross to contribute to the pockets seen in Figs. 4(b) and 4(c). Although the distinction of more appropriate calculation (NM or AFM) is difficult compared to other surface planes, the experimental broad feature at 0.3–0.6 eV along the RAR cut in Fig. 4(f) appears to show a better agreement with the NM calculation in Fig. 4(g). Besides these bulk-originated features, one can find in both Figs. 4(d) and 4(f) several dispersive features within eV of , while they have no counterparts in the calculation. For example, an inner electron band that produces the hexagonal pocket in Fig. 4(b) appears in the (, ) region where the calculated bands in the NM phase are absent. This band displays no band dispersion along the out-of-plane wave vector, [Figs. 4(d) and 4(f)], supporting its surface origin. One can also recognize a hole band producing a small pocket at the X point associated with the pocket in Fig. 4(b). This band is not predicted by the NM calculation and has no dispersion along in the experiment [Fig. 4(g)], and hence, it is also attributed to the SS. Therefore, the (101) surface of RuO2 hosts several dispersive SS that contribute to the formation of surface-derived FS. It is noted here that the inclusion of SOC in the NM calculation slightly improves the matching between the experiment and calculation (see Appendix D), but a fatal disagreement is still found in the AFM calculation even when the SOC is included.
@@ IV Topological surface states
The present study supports the absence of antiferromagnetism and altermagnetic band splitting in bulk single crystal of RuO2 (note that, at the moment, we do not exclude the possibility that a RuO2 thin film shows altermagnetic behavior, because the crystal defects and lattice strain are not the same as those of bulk crystal). Besides this conclusion, we propose that the key characteristics of SS observed at the three different crystal orientations are commonly explained in terms of the non-trivial band topology that is responsible for the emergence of topological SS associated with the bulk DNL. This is enabled by the observation of electronic structures for the (100) and (101) orientations that yields two key findings: (i) the identification of several metallic SSs independent of surface orientations, and (ii) strongly orientation-dependent surface band dispersion (flat vs dispersive). These results are important for understanding the mechanisms underlying spintronic functionalities, because such functionalities are reported by the thin film experiments [19, 33, 5, 6, 63, 65, 22, 66, 9, 20, 28, 42, 73] and thereby the SSs would play a significant role. Importantly, (i) indicates that the emergence of the SSs could be attributed to the bulk-edge correspondence in the topological band structure like the Dirac-cone SS in a strong topological insulator.
To elaborate on this discussion quantitatively, we examine topological nature of the band structure in terms of Berry phase in topological semimetals [10, 13, 17, 12, 25]. Since RuO2 has a rutile structure (space group ) and found to be nonmagnetic from our APRES and first-principles calculations, there exist several nodal lines in the bulk BZ, protected by and mirror symmetries (: space-inversion, : time-reversal; see, Apprndix E). Among these nodal lines, the DNL located near in the mirror plane, referred to here as DNL1 [Fig. 5(a)] is the most important DNL because it is associated with the observed SS. SOC lifts the degeneracy of the DNLs, opening a gap and thereby removing their strict topological protection. However, this does not imply that the original DNLs are topologically trivial. In fact, the Berry curvature generated around the SOC-induced gaps originates from the nontrivial topology of the DNLs that existed before the gap opening, and this curvature can lead to a large intrinsic spin Hall effect [61]. Therefore, to establish the topological nature of RuO2, it is crucial to experimentally verify the existence of surface states that reflect the bulk DNLs, even when the degeneracy is lifted by SOC. Demonstrating such correspondence provides direct evidence that the gapped DNLs inherit their topological character from the underlying band inversion, and thus offers a solid foundation for understanding the spin-current generation mechanism in RuO2. Because the SOC strength in RuO2 is significantly weaker than that in IrO2, this compound provides an ideal platform for investigating the weak-SOC limit where the link between topology and spin transport can be clearly resolved.
In -symmetric material without SOC, the topological invariant describing the DNLs is given by the Zak phase ():
where is a Bloch wave function for th occupied band, and is the wave vector in the surface BZ (for details, see Appendix E) [64, 10, 13]. In the prototypical nodal-line semimetal, the DNL forms a closed loop (i.e. nodal ring), which gives rise to a well-defined region inside the projection of the nodal ring on the surface BZ. Due to the bulk-edge correspondence, this region hosts nearly dispersion-less drumhead-type SSs [10, 13, 17, 12, 25]. In contrast to such Dirac nodal ring, the DNL1 extends across the BZ, connecting its opposite boundaries without forming a closed loop. The DNL1 does not enclose any finite area, and therefore no well-defined “inside” region can be assigned to the phase. Nevertheless, we found that the mirror eigenvalues of bulk bands forming the DNL1 are exchanged along (110)/(10) plane, and this change acts as an effective topological boundary within the BZ from which the SS emerges (for details, see Appendix E). Analogous to the case of nodal line semimetals, each crossing for DNL1 causes a jump in the Zak phase due to the inversion of the occupied-band mirror eigenvalues. This is confirmed by our calculation of Wilson loop encircling a single DNL1 [see the three red circles in Fig. 5(a)] that consistently yields a value of , demonstrating the intrinsic topological nature of DNL1. We evaluated the Wilson loop around three different -loops to examine the nontrivial nature of DNL1 with particular care, to confirm that the Wilson loop, which represents a local Berry phase, indeed captures the global topology of DNL1. Since the two mirror planes, (110) and (10), are symmetry-related and induce opposite phase inversions, the two shifts cancel each other, resulting in a net Zak phase of along the [100] directions. This is manifested in Figs. 5(a) and 5(b) as a blue region in the color map around the projected DNL1, depicted as a wavy cyan line. Thus, no well-defined region can be defined, even though SSs can still emerge around the projections of the nodal lines protected by mirror and symmetries. Namely, the dispersion of SSs of RuO2 on the (100) surface exhibits a drumhead-like shape but lacks a corresponding topological character. Consequently, on the (100) and (110) surfaces where DNL1 is doubly projected, two topological SS (assuming spin-degenerate bands) are expected to emerge from the DNL when SOC is neglected, as illustrated in the left panel of Fig. 5(c). Although SOC opens a finite gap along the former DNL, surface-related states remain in the corresponding -region, reflecting the residual band connectivity that exists without SOC, as shown in the right panel of Fig. 5(c).
To make the above qualitative arguments more quantitative, we carried out slab calculations for the (100) surface. As shown in Fig. 5(f), the slab calculation including SOC predicts two prominent SSs along the cut of surface BZ [for surface BZ, see Fig. 5(d)], both appearing within 0.2 eV of . We confirmed that these two SSs have topological nature, because they merge into the DNL1 along the cut of surface BZ, consistent with the schematics in Fig. 5(c) (for details, see Appendix F). The corresponding ARPES intensity along the same cut (cut 4) shown in Fig. 5(g) reveals one SS (SS2) with strong spectral weight, while the other SS is not clearly resolved, likely because it is largely unoccupied. Systematic ARPES intensity mapping along a few off- cuts (cuts 1–3) allows us to trace the second SS (SS1). Taken together, our ARPES results demonstrate the presence of SSs connecting the bulk DNL, in good agreement with our first-principles calculations, supporting the topologically nontrivial nature of the DNL and associated SSs. We have also carried out the slab calculations on other surface orientations, (110) and (101), and found that the characteristics of the SS and the DNL is universally explained in terms of nontrivial topology and bulk-edge correspondence (Appendix F).
@@
While the present study focuses on the surface of bulk crystal, it would be natural to extend our arguments to the interface between RuO2 (including RuO2 thin film) and other materials. The present study suggests that the outstanding transport properties of RuO2 thin film such as the spin-charge conversion (spin splitting torque), tunneling magnetoresistance, and other physical properties[19, 33, 5, 6, 63, 65, 22, 66, 9, 20, 28, 42, 73, 48]. Such properties involve charge/spin transport across the interface should be discussed with taking into account the existence of prominent topological interface states besides the bulk DNL that strongly contributes to the spin Berry curvature [66, 61]. Our results also show strong crystal-orientation-dependent variation in the band velocity (flat/dispersive) of topological SSs, which would cause significant anisotropy in film properties (for details, see Appendix G). In this regard, an effective model for the surface states and a more elaborate microscopic theoretical framework should be constructed in future studies. It is also worthwhile to comment on the possibility that the catalytic reaction particularly known for the (110) surface [49] may be promoted by the existence of topological SS [61], because high surface conductivity and robustness of the topological SS would be useful for the catalytic reaction [39, 38], working as a “topological catalyst”. In RuO2, the proximity of the topological SS to [ 20 meV ( 200 K) of ; see Fig. 3(f)] for the (110) and (100) surfaces would promote catalytic reaction at room temperature. Hence, we caution that a theoretical model that simply deals with the altermagnetic splitting of bulk bands without incorporating the topological surface/interface states and bulk DNL is insufficient to fully account for some of the experimental results of RuO2.